OpenStudy (anonymous):

PLEASE HELP I got 3/8 is that correct or wrong?? The band at Crest Elementary School has 9 flutes, 6 clarinets, 3 trumpets, 2 trombones, 1 tuba, and 3 French horns. If one instrumentalist is chosen from the band, what is the probability that the student plays the flute?

OpenStudy (anonymous):

I will take a look... one sec

OpenStudy (anonymous):

k thanks :)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

prob = (# flutes) / (total # instruments) = 9 flutes / (9+6+3+2+1+3) = 9 / 24 You can simplify as 3/8 So... You are right!!

OpenStudy (anonymous):

awesome !!!

OpenStudy (anonymous):

if im doing a similar problem and i get 14/50 which is 3.571 what would that simplify to because i really suck at doing that not sure what the process is to do that because i am unable to simply look at that and just KNOW what it is ya know

OpenStudy (anonymous):

7/25

OpenStudy (anonymous):

I usually check to see if both top & bottom are even... if they are, you can always factor out a 2 out of each... just to get started.

OpenStudy (anonymous):

Knowing the decimal answer usually isn't helpful

OpenStudy (anonymous):

oh ok that makes sense

OpenStudy (anonymous):

When you had 9/24, you would need to "just know" that both 9 and 24 are divisible by 3... those multiplication tables you "loved" will come in handy ;)

OpenStudy (anonymous):

k so if its 8/24 is that 1/3?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

keep dividing out anything in common from the top & bottom till you're done and can't do any more. Often it's best to just leave it as a fraction then, not use the calc to get a decimal unless the answer needs to be in a decimal form

OpenStudy (anonymous):

So, if you did probability, 7/25 is fine. Unless you need a probability given as a percentage, then use the calculator

OpenStudy (anonymous):

Georgia built a model of the Empire State Building that is proportional to the actual building. The actual building is 187 feet wide and 1,454 feet tall. Her model is 10 feet tall. Approximately how wide is Georgia's model?

OpenStudy (anonymous):

i did some work and got 1.87 but im not sure if thats right

OpenStudy (anonymous):

187 ---> 10 1454 ---> ? so shrink by multiplying by (10/187) 1454(10/187) = 7.7754 ft

OpenStudy (anonymous):

You can check the shrink factor on the info they give you... 187 (10/187) = 10

OpenStudy (anonymous):

These numbers aren't "pretty" because they used a real life example where the fractions don't easily simplify

OpenStudy (anonymous):

.77?

OpenStudy (anonymous):

What do you mean? This is just like the original question with the poster, except that the numbers are messier. In the original, the scale factor made the finished poster bigger by (5/2). This one makes the model building smaller by (10/187)

OpenStudy (anonymous):

so do i take 187 and divide it by 10

OpenStudy (anonymous):

well that cant be right actually ugh

OpenStudy (anonymous):

yes and no... those 2 numbers are how you make the scaling factor, but you have to be careful to put the right one on top and bottom

OpenStudy (anonymous):

AACK, I messed up

OpenStudy (anonymous):

So sorry!! I had length and width mixed up.

OpenStudy (anonymous):

so i got .053 but the possible answers are A. 0.77 feet B. 2.72 feet C. 1.87 feet D. 1.29 feet

OpenStudy (anonymous):

Shrink that building from 1454 feet tall down to 10 feet tall by multiplying by (10/1454). The width is then 187 in the original down to 187(10/1454) in the model

OpenStudy (anonymous):

1.286 ---> 1.29 (d)

OpenStudy (anonymous):

Uggh... sorry I read it wrong.

OpenStudy (anonymous):

its all good i was confused anyways lol

OpenStudy (anonymous):

original new number ---> number multiply "original" by "(new/original)" to find "new"

OpenStudy (anonymous):

Remember that, and it works for shrink or grow.

OpenStudy (anonymous):

.The actual F-16 has a length of 14.8 meters and a wingspan that measures 9.8 meters. Her model has a length of 0.5 meter. What is the approximate measure of the wingspan on April's model? i got 0.33 is that right?

OpenStudy (anonymous):

14.8 (0.5/14.8) = 0.5 meter for length So for wingspan, 9.8 (0.5/14.8) = 0.33

OpenStudy (anonymous):

In other words, you got it... :)

OpenStudy (anonymous):

see i didnt do it like that at all lol haha

OpenStudy (anonymous):

Miranda hiked to the top of a waterfall. She wanted to know the waterfall's height, but she was unable to measure it. Miranda created a proportion using the information from triangles PQS and TRS to calculate the height of the waterfall. How tall is the waterfall? (Round your answer to the nearest foot.)

OpenStudy (anonymous):

A. 550 ft. B. 806 ft. C. 168 ft. D. 648 ft.

OpenStudy (anonymous):

(looking at the picture now)

OpenStudy (anonymous):

So, triangle STR is similar to triangle SPQ. That means the side ST is proportional to the whole side SP. That proportion is the same proportion as the height TR is to the waterfall height PQ.

OpenStudy (anonymous):

730 (256+730)/(730) = 256 + 730 = 986 For the waterfall, 480 ( 986/730) = 648

OpenStudy (anonymous):

Is that making sense?

OpenStudy (anonymous):

ok yeah it does

OpenStudy (anonymous):

if 8 ft long is converted to 4 inches long then 3 1/2 feet wide would be converted to 1 1/2 inches right?

OpenStudy (anonymous):

8 feet x (4 inches/8 ft) = 4 inches

OpenStudy (anonymous):

probably because the answers are 9 inches 2 inches 1 3/4 inches 1 1/2 inches

OpenStudy (anonymous):

I got 1 3/4 inches the first time I did it. How did you get 1 1/2 ?

OpenStudy (anonymous):

Anyway, looks like it's (c) on your choices

OpenStudy (anonymous):

On that one, you have to be careful to keep the units straight. you can't mix the ft and in directly, so you have to convert something somewhere

OpenStudy (anonymous):

thats why i wasnt getting it haha

OpenStudy (anonymous):

Lots to remember, but you're getting it

OpenStudy (anonymous):

Do you pretty much "get" the idea of scaling up and down now? There may be other ways they try to confuse you or throw you off, but the scaling idea is the same on all

OpenStudy (anonymous):

yeah better its just like every different problem on that that they give me is more and more confusing haha

OpenStudy (anonymous):

what's the next one?

OpenStudy (anonymous):

Max is building a scale model of the solar system. In Max's model, the distance from Earth to the sun is 1.2 feet, and the distance from Venus to the sun is 0.86 feet. If the approximate distance from Earth to the sun is 93 million miles, what is the approximate distance from Venus to the sun?

OpenStudy (anonymous):

Care to set it up and I will check it? Or do you want to walk through it again to be sure you have the concept?

OpenStudy (anonymous):

when i look at this problem i think that i need to take 98 million miles and divide it to 1.2 then that i take that answer and try using that same number to get the answer for the other number but ugh thats probably not right

OpenStudy (anonymous):

Slow down just a step... Say you have a magic wand and can magically make the 1.2 feet distance into 93 million miles. The "magic" is really just multiplying by (93 million / 1.2 )

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

So, Venus starts at 0.86 feet, and "MAGIC" happens... 0.86 (93 million / 1.2) = 66.7 million

OpenStudy (anonymous):

and you "know" which number goes on the bottom, because it's the one you have to get rid of by cancelling in the fraction. So if you start with 1.2, 1.2 must be on the bottom so it cancels to leave 93 million. Now that you have the scale proportion fraction correct, use that "magic" on the other term, so 0.86 inches mult. by the proportion = 67 million

OpenStudy (anonymous):

I think you had this one right too... except you used 98 million instead of 93 million